OpenStudy (anonymous):

What is the general form of the series 1 + 2 + 4 + 8 + 16 + … + 1,024?

rebeccaxhawaii (rebeccaxhawaii):

whats happening to the numbers each time? AND WELCOME TO OS !!!

OpenStudy (mysticfury):

it looks as thought the number are just adding them selves to form the new number

OpenStudy (anonymous):

its a constant addition ?

OpenStudy (anonymous):

the answer choices are so confusing

OpenStudy (mysticfury):

can we see them?

OpenStudy (anonymous):

∑i=1102i−1 ∑i=1102i ∑i=1112i−1∑i=1112i−1 ∑i=1112i

OpenStudy (mysticfury):

*confused face* idk how to make one....... uh .. uh .... uhm... i see what you mean.. they are confusing at least for me

OpenStudy (anonymous):

its it but if it helps the section is called introduction of sequences

OpenStudy (anonymous):

\[\sum_{i=1}^{10}2i-1 thats one of them

OpenStudy (anonymous):

then the 2i-1 becomes 2 and the others the 10 becomes 11 and 11 1i-1